# Yang-Lee Zeros in Quantum Phase Transition: An Entanglement Perspective

### v1

@@ Abstract
We extend the Yang-Lee theory to quantum phase transitions to show how singularity enters the phase transition points in one-dimensional many-body systems. We primarily focus on the distribution of Yang-Lee zeros and its associated Yang-Lee edge singularity of two prototypical models: the Su-Schrieffer-Heeger (SSH) model and the XXZ spin chain. By taking the zero-temperature limit, we show how the Yang-Lee zeros approach the quantum phase transition points on the complex plane of parameters. To characterize the edge singularity induced by Yang-Lee zeros in quantum phase transition, we introduce the entanglement entropy of the ground state to show the edges of Yang-Lee zeros lead to the entanglement transition. We further show our results are also applicable to the general non-interacting parity-time-symmetric Hamiltonians.
@@
Introduction.—Yang-Lee zeros Yang and Lee (1952); Lee and Yang (1952), defined as the zero points of the partition function of the canonical ensemble, give a mathematical explanation of the nonanalyticity of thermodynamic observables at phase transition points. In the classical ferromagnetic Ising model, Yang and Lee investigated the zeros of the partition function in the presence of an imaginary magnetic field in order to understand the origin of the singularity in the ferromagnetic phase transition when one increases the temperature. The thermal phase transition takes place when Yang-Lee zeros touch the positive real axis on the complex plane of fugacity in the thermodynamic limit. In the vicinity of the edge of the distribution of Yang-Lee zeros, critical phenomena collectively known as Yang-Lee edge singularity emerge, which are accompanied by anomalous scaling laws Fisher (1978); Kurtze and Fisher (1979); Fisher (1980); Cardy (1985); Cardy and Mussardo (1989); Zamolodchikov (1991); Bena et al. (2005); Fisher (1965). Yang-Lee theory is applicable to a variety of thermal phase transitions in classical Biskup et al. (2000); Arndt (2000); Blythe and Evans (2002); Lee (2013) and quantum von Gehlen (1991); Sumaryada and Volya (2007); Abraham et al. (1996); Matsumoto et al. (2022); Vecsei et al. (2022); Gnatenko et al. (2017); Peotta et al. (2021); Shen et al. (2023); BENA et al. (2005) systems.
@@
Despite the great success of the application of Yang-Lee zeros to explain the origin of the singularity in thermal phase transitions, how to apply Yang-Lee zeros to understand quantum phase transitions remains largely unexplored Jian et al. (2021); Kist et al. (2021); Li et al. (2023). The singularity of a quantum phase transition arises at absolute zero when we manipulate the parameters in a Hamiltonian and can be universally characterized by the singular behavior of the entanglement pattern in ground-state wavefunctions Levin and Wen (2006); Kitaev and Preskill (2006); Fidkowski (2010). Therefore, it is expected to universally characterize the Yang-Lee edge singularity in terms of the singularity in quantum entanglement.
@@
In this Letter, we develop a Yang-Lee theory for quantum phase transitions and focus on how the Yang-Lee zeros approach the phase transition points by taking the zero-temperature limit. First, we investigate the distribution of Yang-Lee zeros in the parameter space of a non-interacting model: the SSH model Chang et al. (2020) by generalizing the Rice-Mele coupling Lin et al. (2020) to an imaginary one. We explore the nonunitary criticality of the non-Hermitian SSH model originating from Yang-Lee edge singularity. These quantum critical phenomena are induced by the exceptional points where a nonanalytic excitation spectrum exists. Furthermore, we figure out that the edges of the distribution of Yang-Lee zeros at absolute zero correspond to the entanglement transition points of the ground states. We also generalize our discussion to general parity-time (PT)-symmetric free-fermion systems to unveil hitherto unnoticed universality in Yang-Lee edge singularity in terms of transitions in the entanglement entropy.
@@
We also discuss the Yang-Lee theory of an interacting model: the XXZ spin chain. On the complex plane of the anisotropy, we investigate the distribution of Yang-Lee zeros near the ferromagnetic phase transition point. By analyzing the Bethe ansatz of the XXZ model, we prove that the Yang-Lee zeros approach the phase transition point in a line for arbitrary system size. The edge of the distribution of Yang-Lee zeros turns out to be the ferromagnetic phase transition point. We also show the edge of the Yang-Lee zeros corresponds to an entanglement transition. This exact result paves the way for figuring out Yang-Lee zeros in quantum phase transitions of interacting models. Finally, we provide a general argument for the correspondence between Yang-Lee edge singularity and entanglement transition.
@@
Non-interacting Model.—We begin from a non-interacting model: the non-Hermitian SSH model. The Hamiltonian in the momentum space is defined as Lieu (2018); Chang et al. (2020)
@@
where $k$ is the momentum and $c_{kA}^{(\dagger)}$ and $c_{kB}^{(\dagger)}$ represent the annihilation (creation) operators of the fermions with momentum $k$ on the sublattice $A$ and $B$ respectively. The energy spectrum of the Bloch Hamiltonian is $E_{k}^{\pm}=\pm\sqrt{|v+we^{ik}|^{2}-u^{2}}=:\pm E_{k}$. The model has three phases: the trivial PT-unbroken gapped phase for $w-v<-u$, the PT-broken gapless phase for $-u<w-v<u$, and the topologically nontrivial PT-unbroken gapped phase for $w-v>u$ Chang et al. (2020).
@@
$\text{Re}[\beta E_{k}]=0,\text{Im}[\beta E_{k}]=(2n+1)\pi/\beta$ (4)
with $n\in\mathbb{Z}$. In the zero-temperature limit $\beta\to\infty$, the conditions are satisfied in the thermodynamic limit for the gapless phase $|v-w|\leq u$, corresponding to the whole PT-broken phase. In the gapless phase, there always exists an interval in the momentum space with a finite purely imaginary part of the energy spectrum, which induces the Yang-Lee zeros. In Fig. 1, we show how Yang-Lee zeros approach the quantum phase transition points $|w-v|=u$ by taking the zero-temperature limit. To further understand the distribution of Yang-Lee zeros, we introduce the number of roots of the partition function $\chi$ Li et al. (2023), which is defined as the number of the integer $n$ satisfying the condition of zeros in Eq. (4). In the zero-temperature limit, the number of roots $\chi$ takes the form of
@@
where $u>|v-w|$ in the PT-broken phase. Near the phase boundary $|v-w|=u$, the asymptotic behavior of $\chi$ can be shown as $\chi\propto(u-|v-w|)^{1/2}$ which is attributed to the exceptional points Li et al. (2023) in the gapless phase. Since the dispersion relation is $E_{k}^{\pm}=\pm\sqrt{v^{2}+w^{2}+2vw\cos k-u^{2}}$, the gap closes at
@@
with $\delta=|v-w|-u$ and $\xi=vw/(2u\delta)$ is the correlation length. See Supplemental Material Sup for details. Hence, from Eq. (8), we obtain the critical exponent $\nu=1$ which is defined as $\xi^{-1}\propto\delta^{\nu}$ Sachdev (2011). In addition, from the correlation function (8), we have the anomalous scaling dimension $\eta=3/2$ defined as $\lim_{x\to\infty}C(x)\propto\exp(-x/\xi)/x^{d-2+\eta}$ Sachdev (2011). The anomalous power $1/2$ originates from the dispersion relation $E\propto(k-k_{E})^{1/2}$ near the exceptional points. Hence, combining with the dynamical critical exponent $z=1$, we have the critical exponents for the Yang-Lee edge singularity here, which are similar to the ones in Yang-Lee singularity in the non-Hermitian BCS superconductivity Li et al. (2023). They are both induced by exceptional points in a PT-symmetric Hamiltonian.
Now we turn to discuss the singularity induced by Yang-Lee zeros in terms of the entanglement entropy (EE) of the ground state. The EE of the subsystem $A$ is defined as $S_{A}:=\text{Tr}_{A}(\rho_{A}\log\rho_{A})$ where $\rho_{A}=\text{Tr}_{\bar{A}}|G_{R}\rangle\langle G_{L}|$ Chang et al. (2020). The EE discontinuously jumps from a subarea-law scaling $S_{A}=(1/6)\log L_{A}$ in the PT-broken gapless phase, where Yang-Lee zeros are located, to an area-law scaling $S=\text{const}$ in the PT-unbroken gapped phase Rudner and Levitov (2009); Chang et al. (2020). Here $L_{A}$ denotes the length of the subsystem $A$. We can see the edges of the distribution of zeros correspond to the entanglement transitions. We note that the correspondence is also applicable to the Hermitian SSH model. In the Hermitian case $u=0$, the system at the phase transition point $v=w$, which is the edge of Yang-Lee zeros, exhibits a subarea-law scaling for the ground-state EE. To elucidate the generality of the correspondence, we consider an arbitrary local PT-symmetric free-fermion Hamiltonian. By generalizing the discussion in Ref. Wolf (2006), we can prove that the gapless PT-broken phase with Yang-Lee zeros shows at least a subarea-law scaling $S_{A}\geq c_{-}\log L_{A}$ with $c_{-}>0$ and the gapped PT-unbroken phase shows an area-law scaling. Therefore, the correspondence between the edges of distribution of Yang-Lee zeros and the entanglement transitions is general for PT-symmetric free-fermion Hamiltonians. The Yang-Lee zeros introduce the singularity to the entanglement transition. In addition, the correspondence can be also seen in the transverse-field XY model (See Supplemental Material Sup ).
Interacting Model.— Above we have considered the Yang-Lee zeros in the quantum phase transitions of the non-interacting cases. In this part, we extend our discussion to an interacting case. We investigate the XXZ model to understand the emergence of Yang-Lee zeros from the Bethe ansatz. The Hamiltonian of the XXZ spin chain is defined as
@@
where $J>0$, $L$ is the size of the spin chain and we extend the anisotropy parameter $\Delta$ to a complex one to consider the distribution of Yang-Lee zeros. We here focus on the ferromagnetic phase transition point $\Delta=1$. In the Hermitian case, the ground state of the gapped (gapless) phase $\Delta>1$ ($-1<\Delta<1$) is (anti)ferromagnetic Takahashi (2005).
To consider the condition of Yang-Lee zeros, we follow the definition of the partition function in Eq. (Yang-Lee Zeros in Quantum Phase Transition: An Entanglement Perspective): $Z=\sum_{n}{}_{L}\langle E_{n}|\exp(-\beta H)|E_{n}\rangle_{R}$. We show a theorem:
@@
From Eq. (10), we deduce that there are overall $N$ complex roots of $z$ which can be defined as $z_{1},\cdots,z_{N}$, where $N=(L/2)^{2}$ for even $L$ and $N=(L^{2}-1)/4$ for odd $L$. By rewriting Eq. (10) as $\sum_{M=1}^{N}a_{M}z^{M}=0$, we have $a_{0}=2$ and $a_{N}=1$ for an even $L$ or $a_{0}=a_{N}=2$ for an odd $L$, indicating that $z=0$ and $z=\infty$ are both not the solutions of Eq. (10). Therefore, all the roots $z$ are finite that gives rise to the zeros $\Delta_{j}$ on the complex plane of $\Delta$ as
@@
where $n\in\mathbb{Z}$ and $j=1,\cdots,N$. We can see the zeros form a line around the phase transition point. In the zero-temperature limit, those zeros are equivalent to $\Delta_{j}=1+i(L-1)\frac{2n\pi}{\beta J}$ which are nothing but the line $\text{Re}\Delta=1$. This indicates that the Yang-Lee zeros approach the real axis of $\Delta$ through the line $\text{Re}\Delta=1$. Below we outline the proof of Theorem 1.
@@
To begin with, we explain the structure of the exact solution of the XXZ spin chain. The Bethe ansatz equations (BAE) of the XXZ chain are given by Yang and Yang (1966a, b); Takahashi (2005)
@@
where we define $x_{j}$ as $\exp(ik_{j})=\sin\frac{\phi}{2}(x_{j}+i)/\sin\frac{\phi}{2}(x_{j}-i)$ and $\phi:=\cosh^{-1}\Delta$ with $k_{j}$ being the quasimomenta of the excited magnons. In Eq. (12), $j=1,\cdots,M$ where $M$ represents the number of magnons. The energy of the $M$-magnon sector takes the form of Takahashi (2005)
@@
where $E_{0}=-JL\Delta/4$ represents the ground state energy for the states with all spins up or down. Near the ferromagnetic phase transition point $\Delta=1$, we introduce a quantity $\delta:=\Delta-1$ and the quantity $\phi$ can be approximated as $\phi\approx\sqrt{2\delta}$. To understand the behavior of Yang-Lee zeros, we first turn to a limit $\delta\to 0$, under which the system becomes the Heisenberg model. We can see the $(L+1)$-degeneracy of the ground state in the Heisenberg model. Those ground states can be constructed as $|M\rangle=(S^{-})^{M}|\Omega\rangle$ where we denote $|\Omega\rangle$ as the all-spin-up state. However, the $(L+1)$-degeneracy will be lifted if we add a perturbation to the parameter $\Delta$. To analyze the gap between levels, we consider the expansion of the BAE for XXZ chain (12) around $\phi=0$:
@@
which we call simplified Bethe ansatz equations (SBAE) and are shown to be a set of algebraic equations of $\cot(\phi x_{j}/2)$. From the equations one can prove that the solutions $\cot(\phi x_{j}/2)$ only depend on the size $L$. Meanwhile, the energy can be also expressed as a function of $\cot(\phi x_{j}/2)$:
@@
By substituting the solutions of the SBAE into Eq. (15), we obtain the energy of the $M$-magnon sector as
@@
which is equivalent to Eq. (10). The maximal value of the exponent $M(L-M)$ is $L^{2}/4$ for an even $L$ with $M=L/2$ or $(L^{2}-1)/4$ for an odd $L$ with $M=(L-1)/2,(L+1)/2$. This completes the proof. Details of the proof are attached in the Supplemental Material Sup . ∎
@@
Figure 2: (a): The distribution line of Yang-Lee zeros in the non-Hermitian XXZ model on the complex plane of $\Delta$. The size of the spin chain is $L=6$ and the inverse temperature is $\beta=100$. The zeros of the partition function approach the phase transition point $\Delta=1$ through a straight line. The dashed lines represent the entanglement between spins. Here we show the nearest 200 zeros to the phase transition point. (b): The scaling of EE $S_{A}$ of subsystem $A$ with size $L_{A}$ at the point $\Delta=0.99+0.01i$. The fitting curve is chosen as $S_{A}=1.881+0.184\log(\sin(\pi L_{A}/L)))$, exhibiting the subarea law.
We can further understand the line distribution of Yang-Lee zeros from the gap behavior of the non-Hermitian XXZ model on the complex plane. For the Hamiltonians with $\text{Re}\Delta>1$, we can see the systems must be gapped from Eq. (16), with the gap $\Delta E=J\text{Re}\delta$. However, for those Hamiltonians with $\text{Re}\Delta<1$, the systems become gapless. In this phase, the ground state has $M=L/2$ magnons and the real part of the energy becomes $\text{Re}E=E_{0}+\frac{(L/2)^{2}}{L-1}J\text{Re}\delta$ with $\text{Re}\delta<0$ when $L$ is even. Then the energy level distance between the ground state and the first excited state is given by
@@
Next, we examine the critical exponents for the Yang-Lee edge singularity of the non-Hermitian XXZ model. We introduce a magnetic field $h$ to the Hamiltonian. By solving the BAE (12), the susceptibility in the gapless phase takes the form of $\lim_{h\to 0}\chi=\frac{2\gamma}{J\pi\delta}$ near the critical point where $\gamma$ is a constant Takahashi (2005). Therefore, we obtain the critical exponent $\sigma=1$ from $\chi$ defined as $\chi\sim\delta^{-\sigma}$. At the critical point, the susceptibility diverges. In addition, By definition of the density of Yang-Lee zeros $g(\Delta)=\sum_{j}\delta(\Delta-\Delta_{j})$, we have
@@
Furthermore, here the edge of the Yang-Lee zeros is also associated with the entanglement transition of the ground state. The zeros are distributed on the line $\text{Re}\Delta=1$ and approach the phase transition point $\Delta=1$ when we decrease the temperature. Hence, we can consider the phase transition point as the edge of the Yang-Lee zeros. On the real axis, the EE of the gapless region $\text{Re}\Delta<1$ shows a subarea-law scaling while that of the gapped region $\text{Re}\Delta>1$ shows an area-law scaling. Therefore, the correspondence between the Yang-Lee edge singularity and the entanglement transitions is applicable to not only the non-interacting cases but also the interacting systems. We note that the points on the distribution line of the Yang-Lee zeros also represent the entanglement transition points. To illustrate this, we calculate the EE of the model at $\Delta=0.99+0.01i$ with non-Hermitian Density-Matrix Renormalization Group Yamamoto et al. (2022); Fishman et al. (2022) in Fig. 2(b), which also shows a subarea-law scaling. The reduced density matrix of the subsystem $A$ is defined as $\rho_{A}=\text{Tr}_{\bar{A}}|\psi_{0}^{R}\rangle\langle\psi_{0}^{R}|$, where $|\psi_{0}^{R}\rangle$ is the right eigenstate of the Hamiltonian (9) with the lowest real energy. On the other hand, the systems with $\text{Re}\Delta>1$ show the area law in the gapped phase since the ground state in this phase is a direct-product state. Therefore, all of the Yang-Lee zeros correspond to entanglement transition points on the complex plane.

### v2

@@ Abstract
We study the Yang-Lee theory in quantum phase transitions from the perspective of quantum entanglement in one-dimensional many-body systems. We primarily focus on the distribution of Yang-Lee zeros and its associated Yang-Lee edge singularity of two prototypical models: the Su-Schrieffer-Heeger model and the XXZ spin chain. By taking the zero-temperature limit, we show how the Yang-Lee zeros approach the quantum phase transition points on the complex plane of parameters. To characterize the edge singularity induced by Yang-Lee zeros in quantum phase transition, we introduce the entanglement entropy of the ground state to show that the edges of Yang-Lee zeros lead to the ground-state entanglement transition. We further show that our results are also applicable to the general non-interacting parity-time-symmetric Hamiltonians.
I Introduction
@@
Yang-Lee zeros [1, 2], defined as the zero points of the partition function of the canonical ensemble, give a mathematical explanation of the nonanalyticity of thermodynamic observables at phase transition points. In the classical ferromagnetic Ising model, Yang and Lee investigated the zeros of the partition function in the presence of an imaginary magnetic field in order to understand the origin of the singularity in the ferromagnetic phase transition when one increases the temperature. The thermal phase transition takes place when Yang-Lee zeros touch the positive real axis on the complex plane of fugacity in the thermodynamic limit. In the vicinity of the edge of the distribution of Yang-Lee zeros, critical phenomena collectively known as Yang-Lee edge singularity emerge, which are accompanied by anomalous scaling laws [3, 4, 5, 6, 7, 8, 9, 10, 11]. Yang-Lee theory is applicable to a variety of thermal phase transitions in classical [12, 13, 14, 15] and quantum [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28] systems.
@@
Despite the great success of the application of Yang-Lee zeros to explain the origin of the singularity in thermal phase transitions, there exist exceptions in quantum phase transitions such as the topological quantum phase transition [29, 30], which are characterized by the change of the topological invariants instead of singularity in thermodynamic quantities. With the introduction of quantum entanglement, it was discovered that the singularity in a quantum phase transition can be more universally characterized by the singular behavior of the entanglement pattern in ground-state wavefunctions [31, 32, 33, 34]. Therefore, it is expected to universally characterize the Yang-Lee edge singularity in terms of the singularity in quantum entanglement instead of thermodynamic observables in the traditional Yang-Lee theory.
@@
In this work we develop a Yang-Lee theory for quantum phase transitions and focus on how the Yang-Lee zeros approach the phase transition points by taking the zero-temperature limit. First, we investigate the distribution of Yang-Lee zeros in the parameter space of a noninteracting model: the Su-Schrieffer-Heeger (SSH) model [35] by generalizing the Rice-Mele coupling [36] to an imaginary one. We explore the nonunitary criticality of the non-Hermitian SSH model originating from Yang-Lee edge singularity. These quantum critical phenomena are induced by the exceptional points where a nonanalytic excitation spectrum exists. Furthermore, we figure out that the edges of Yang-Lee zeros at absolute zero correspond to the entanglement transition points of the ground states. We also generalize our discussion to general parity-time (PT)-symmetric free-fermion systems to unveil universally unnoticed universality in Yang-Lee edge singularity in terms of transitions in the entanglement entropy. This result indicates that entanglement entropy may serve as a good characterization for the universality class of Yang-Lee edge singularity in quantum phase transitions. In contrast to Ref. [34], we have rigorously proved the correspondence between the edges of Yang-Lee zeros and ground-state entanglement transitions for all PT-symmetric free-fermion systems rather than entanglement transitions of the long-time steady states of a class of specific models.
@@
We also discuss the Yang-Lee theory of an interacting model: the XXZ spin chain. On the complex plane of the anisotropy, we investigate the distribution of Yang-Lee zeros near the ferromagnetic phase transition point. By analyzing the Bethe ansatz of the XXZ model, we prove that the Yang-Lee zeros approach the phase transition point in a line for arbitrary system size. The edge of the distribution of Yang-Lee zeros turns out to be the ferromagnetic phase transition point. We also show that the edge of the Yang-Lee zeros corresponds to an entanglement transition. This exact result paves the way for figuring out Yang-Lee zeros in quantum phase transitions of interacting models. Finally, we provide a general argument for the correspondence between Yang-Lee edge singularity and entanglement transition.
II Non-interacting Model
@@
We begin from a non-interacting model: the non-Hermitian SSH model. The Hamiltonian in the momentum space is defined as [37, 35]
@@
where $k$ is the momentum and $c_{kA}^{(\dagger)}$ and $c_{kB}^{(\dagger)}$ represent the annihilation (creation) operators of the fermions with momentum $k$ on the sublattice $A$ and $B$ respectively. The energy spectrum of the Bloch Hamiltonian is $E_{k}^{\pm}=\pm\sqrt{|v+we^{ik}|^{2}-u^{2}}=:\pm E_{k}$. The model has three phases: the trivial PT-unbroken gapped phase for $w-v<-u$, the PT-broken gapless phase for $-u<w-v<u$, and the topologically nontrivial PT-unbroken gapped phase for $w-v>u$ [35].
@@
$\text{Re}[E_{k}]=0,\text{Im}[E_{k}]=(2n+1)\pi/\beta$ (4)
with $n\in\mathbb{Z}$. To figure out the distribution of Yang-Lee zeros, we first assume a finite $\beta$ and investigate where the condition (4) is satisfied in the momentum space when we take the thermodynamic limit. Then we take the zero-temperature limit to see how the distribution of Yang-Lee zeros approaches the quantum phase transition points. The order of taking the two limits cannot be exchanged. In the zero-temperature limit $\beta\to\infty$, we find that the condition (4) is satisfied in the thermodynamic limit for the gapless phase $|v-w|\leq u$, corresponding to the whole PT-broken phase. In the gapless phase, there always exists an interval in the momentum space with a finite purely imaginary part of the energy spectrum, which induces the Yang-Lee zeros. In Fig. 1, we show how Yang-Lee zeros approach the quantum phase transition points $|w-v|=u$ by taking the zero-temperature limit. To further understand the distribution of Yang-Lee zeros, we introduce the number of roots of the partition function $\chi$ [26], which is defined as the number of the integer $n$ satisfying the condition of zeros in Eq. (4). The number of roots $\chi$ can be understood as the density of Yang-Lee zeros on the plane of parameters. In the zero-temperature limit, the number of roots $\chi$ takes the form of
@@
where $u>|v-w|$ in the PT-broken phase. Near the phase boundary $|v-w|=u$, the asymptotic behavior of $\chi$ can be shown as $\chi\propto(u-|v-w|)^{1/2}$ which is attributed to the exceptional points [26] in the gapless phase. Since the dispersion relation is $E_{k}^{\pm}=\pm\sqrt{v^{2}+w^{2}+2vw\cos k-u^{2}}$, the gap closes at
@@
where $\delta=|v-w|-u$ and $\xi=vw/(2u\delta)$ is the correlation length. See Appendix A for details. Hence, from Eq. (8), we obtain the critical exponent $\nu=1$ which is defined as $\xi^{-1}\propto\delta^{\nu}$ [38]. In addition, from the correlation function (8), we have the anomalous scaling dimension $\eta=3/2$ defined as $\lim_{x\to\infty}C(x)\propto\exp(-x/\xi)/x^{d-2+\eta}$ [38]. The anomalous power $1/2$ originates from the dispersion relation $E\propto(k-k_{E})^{1/2}$ near the exceptional points. Hence, combining with the dynamical critical exponent $z=1$, we have the critical exponents for the Yang-Lee universality class here, which are similar to the ones in the Yang-Lee universality class in the non-Hermitian BCS superconductivity [26]. They are both induced by exceptional points in a PT-symmetric Hamiltonian.
Now we turn to discuss the singularity induced by Yang-Lee zeros in terms of the entanglement entropy (EE) of the ground state. The EE of the subsystem $A$ is defined as $S_{A}:=\text{Tr}_{A}(\rho_{A}\ln\rho_{A})$, where $\rho_{A}=\text{Tr}_{\bar{A}}|G_{R}\rangle\langle G_{L}|$ [35]. The EE discontinuously jumps from a subarea-law scaling $S_{A}=(1/6)\ln L_{A}$ in the PT-broken gapless phase, where Yang-Lee zeros are located, to an area-law scaling $S=\text{const}$ in the PT-unbroken gapped phase [39, 35]. Here $L_{A}$ denotes the length of the subsystem $A$. We can see that the edges of the distribution of zeros correspond to the entanglement transitions. We note that the correspondence is also applicable to the Hermitian SSH model. In the Hermitian case $u=0$, the system at the phase transition point $v=w$, which is the edge of Yang-Lee zeros, exhibits a subarea-law scaling for the ground-state EE. To elucidate the generality of the correspondence, we consider an arbitrary local PT-symmetric free-fermion Hamiltonian. By generalizing the discussion in Ref. [40], we prove that the gapless PT-broken phase with Yang-Lee zeros shows at least a subarea-law scaling $S_{A}\geq c_{-}\ln L_{A}$ with $c_{-}>0$ and the gapped PT-unbroken phase shows an area-law scaling in the one-dimensional case. Therefore, the correspondence between the edges of the distribution of Yang-Lee zeros and the entanglement transitions is general for PT-symmetric free-fermion Hamiltonians. The Yang-Lee zeros introduce the singularity to the entanglement transition. Thus, entanglement entropy can be considered as a universal characterization of the Yang-Lee edge singularity in quantum phase transitions. See Appendix. B for the details of the proof.
III Interacting Model
Above we have considered the Yang-Lee zeros in the quantum phase transitions of the non-interacting cases. In this part, we extend our discussion to an interacting case. We investigate the XXZ model to understand the emergence of Yang-Lee zeros from the Bethe ansatz. The Hamiltonian of the XXZ spin chain is defined as
@@ III Interacting Model
where $J>0$, $L$ is the size of the spin chain and we extend the anisotropy parameter $\Delta$ to a complex one to consider the distribution of Yang-Lee zeros. Here we focus on the ferromagnetic phase transition point $\Delta=1$. In the Hermitian case, the ground state of the gapped (gapless) phase $\Delta>1$ ($-1<\Delta<1$) is (anti)ferromagnetic [41].
To consider the condition of Yang-Lee zeros, we follow the definition of the partition function in Eq. (II): $Z=\sum_{n}{}_{L}\langle E_{n}|\exp(-\beta H)|E_{n}\rangle_{R}$. We show the following theorem:
@@
From Eq. (10), we deduce that overall, there are $N$ complex roots of $z$ which can be defined as $z_{1},\cdots,z_{N}$, where $N=(L/2)^{2}$ for even $L$ and $N=(L^{2}-1)/4$ for odd $L$. By rewriting Eq. (10) as $\sum_{M=1}^{N}a_{M}z^{M}=0$, we have $a_{0}=2$ and $a_{N}=1$ for an even $L$ or $a_{0}=a_{N}=2$ for an odd $L$, indicating that both $z=0$ and $z=\infty$ are not the solutions of Eq. (10). Therefore, all the roots $z$ are finite that give rise to the zeros $\Delta_{j}$ on the complex plane of $\Delta$ as
@@
where $n\in\mathbb{Z}$ and $j=1,\cdots,N$. We can see that the zeros form a line around the phase transition point. In the zero-temperature limit, those zeros are equivalent to $\Delta_{j}=1+i(L-1)\frac{2n\pi}{\beta J}$, which are simply the line $\text{Re}\Delta=1$. This indicates that the Yang-Lee zeros approach the real axis of $\Delta$ through the line $\text{Re}\Delta=1$. Below we outline the proof of Theorem 1.
@@
To begin with, we explain the structure of the exact solution of the XXZ spin chain. The Bethe ansatz equations (BAEs) of the XXZ chain are given by [42, 43, 41]
@@
where we define $x_{j}$ as $\exp(ik_{j})=\sin\frac{\phi}{2}(x_{j}+i)/\sin\frac{\phi}{2}(x_{j}-i)$ and $\phi:=\cosh^{-1}\Delta$ with $k_{j}$ being the quasimomenta of the excited magnons. In Eq. (12), $j=1,\cdots,M$ where $M$ represents the number of magnons. The energy of the $M$-magnon sector takes the form of [41]
@@
where $E_{0}=-JL\Delta/4$ represents the ground state energy for the states with all spins up or down. Near the ferromagnetic phase transition point $\Delta=1$, we introduce a quantity $\delta:=\Delta-1$ and the quantity $\phi$ can be approximated as $\phi\approx\sqrt{2\delta}$. To understand the behavior of Yang-Lee zeros, we first turn to a limit $\delta\to 0$, under which the system becomes the Heisenberg model. We can see the $(L+1)$-degeneracy of the ground state in the Heisenberg model. Those ground states can be constructed as $|M\rangle=(S^{-})^{M}|\Omega\rangle$ where we denote $|\Omega\rangle$ as the all-spin-up state. However, the $(L+1)$-degeneracy will be lifted if we add a perturbation to the parameter $\Delta$. To analyze the gap between levels, we consider the expansion of the BAEs for XXZ chain (12) around $\phi=0$,
@@
which we call simplified Bethe ansatz equations (SBAEs) and are shown to be a set of algebraic equations of $\cot(\phi x_{j}/2)$. From the equations, one can prove that the solutions $\cot(\phi x_{j}/2)$ only depend on the size $L$. Meanwhile, the energy can also be expressed as a function of $\cot(\phi x_{j}/2)$,
@@
By substituting the solutions of the SBAEs into Eq. (15), we obtain the energy of the $M$-magnon sector as
@@
which is equivalent to Eq. (10). The maximal value of the exponent $M(L-M)$ is $L^{2}/4$ for an even $L$ with $M=L/2$ or $(L^{2}-1)/4$ for an odd $L$ with $M=(L-1)/2,(L+1)/2$. This completes the proof. Details of the proof are given in Appendix. C
@@
Figure 2: (a): The distribution line of Yang-Lee zeros in the non-Hermitian XXZ model on the complex plane of $\Delta$. The size of the spin chain is $L=6$ and the inverse temperature is $\beta=100$. The zeros of the partition function approach the phase transition point $\Delta=1$ through a straight line. The dashed lines represent the entanglement between spins. Here we show the nearest 200 zeros to the phase transition point. (b): The scaling of EE $S_{A}$ of subsystem $A$ with size $L_{A}$ at the point $\Delta=0.99+0.01i$ where $L=200$. The fitting curve is chosen as $S_{A}=1.881+0.184\log(\sin(\pi L_{A}/L)))$, exhibiting the subarea law.
We can further understand the line distribution of Yang-Lee zeros from the gap behavior of the non-Hermitian XXZ model on the complex plane. For the Hamiltonians with $\text{Re}\Delta>1$, we can see that the systems must be gapped from Eq. (16), with the gap $\Delta E=J\text{Re}\delta$. However, for those Hamiltonians with $\text{Re}\Delta<1$, the systems become gapless. In this phase, the ground state has $M=L/2$ magnons and the real part of the energy becomes $\text{Re}E=E_{0}+\frac{(L/2)^{2}}{L-1}J\text{Re}\delta$ with $\text{Re}\delta<0$ when $L$ is even. Then the energy level distance between the ground state and the first excited state is given by
@@
Next, we examine the critical exponents for the Yang-Lee edge singularity of the non-Hermitian XXZ model. We introduce a magnetic field $h$ to the Hamiltonian. By solving the BAE (12), the susceptibility in the gapless phase takes the form of $\lim_{h\to 0}\chi=\frac{2\gamma}{J\pi\delta}$ near the critical point where $\gamma$ is a constant [41]. Therefore, we obtain the critical exponent $\sigma=1$ from $\chi$ defined as $\chi\sim\delta^{-\sigma}$. At the critical point, the susceptibility diverges. In addition, by definition of the density of Yang-Lee zeros $g(\Delta)=\sum_{j}\delta(\Delta-\Delta_{j})$, we have
@@
Furthermore, here the edge of the Yang-Lee zeros is also associated with the entanglement transition of the ground state. The zeros are distributed on the line $\text{Re}\Delta=1$ and approach the phase transition point $\Delta=1$ when we decrease the temperature. Hence, we can consider the phase transition point as the edge of the Yang-Lee zeros. On the real axis, the EE of the gapless region $\text{Re}\Delta<1$ shows a subarea-law scaling while that of the gapped region $\text{Re}\Delta>1$ shows an area-law scaling. Therefore, the correspondence between the Yang-Lee edge singularity and the entanglement transitions is applicable to not only the non-interacting cases but also the interacting systems. We note that the points on the distribution line of the Yang-Lee zeros also represent the entanglement transition points. To illustrate this, we calculate the EE of the model at $\Delta=0.99+0.01i$ with the non-Hermitian density-matrix renormalization group [44, 45] in Fig. 2(b), which also shows a subarea-law scaling. The reduced density matrix of the subsystem $A$ is defined as $\rho_{A}=\text{Tr}_{\bar{A}}|\psi_{0}^{R}\rangle\langle\psi_{0}^{R}|$, where $|\psi_{0}^{R}\rangle$ is the right eigenstate of the Hamiltonian (9) with the lowest real energy. On the other hand, the systems with $\text{Re}\Delta>1$ show the area law in the gapped phase since the ground state in this phase is a direct-product state. Therefore, all of the Yang-Lee zeros correspond to entanglement transition points on the complex plane.